OpenStudy (anonymous):

1 and 2 are a linear pair. m1 1= x - 30, and M2 = x+76 find the measure of each angle

OpenStudy (mazefordays090):

i have no idea

OpenStudy (anonymous):

exactly >_>

OpenStudy (mazefordays090):

lol

OpenStudy (anonymous):

Well I've been stuck working on this for the past hour It's a multiple choice question <1= 67 <2 = 123 <1=37 <2 = 153 <1=37 <2 = 143 <1=67 <2 = 113

OpenStudy (anonymous):

m1 = x - 30, and m2 = x+76

OpenStudy (anonymous):

m1 = x - 30, and m2 = x+76 when two measuments are a 'linear pair' what that means is that when you add the two measured angles together they are suppose to equal 180 degrees. like this: m1 + m2 = 180 so, x-30 + x+76 = 180 combine like terms 2x + 46 = 180 isolate x 2x = 180-46 2x = 137 x = 137/2 x = 67 back substitute in x into the original measured angles given M1 = x-30 = 67-30 = 37 now, back substitute in x for M2 and you will have the other half of the question done ^_^

OpenStudy (anonymous):

I gotta take off, I hope this helped! ^_^

OpenStudy (anonymous):

Thanks!